OpenStudy (anonymous):

Evaluate the following expressions when a = -2, b = 3, c = -4 and d = -6. You can use the backslash key to indicate absolute value. \-5\ = 5 You may also have a | symbol that can be accessed by pressing the Shift key and the \ key at the same time. |-5|=5. 1.|a+b| 2.5|c+b| 3.a+b|c| 4.|a+c|/-d 5. 3|a+d|+b

OpenStudy (jack1):

1 a+b =...?

OpenStudy (jack1):

@andrewj43 1. a=-2 and b= 3 so a +b =...?

OpenStudy (anonymous):

i have to show the steps

OpenStudy (jack1):

these are the steps: Q1: a+b =...?

OpenStudy (jack1):

dude, you there?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

sweet, Q1: a+b =...? if a = -2 and b = +3

OpenStudy (anonymous):

it would be 2 +3

OpenStudy (jack1):

no, -2 +3, the absolute value part happens at the end once the equation inside the lines is complete

OpenStudy (jack1):

if it asked |a| + |b| then it would be 2+3

OpenStudy (anonymous):

oh so then it would be 1

OpenStudy (jack1):

yep, and its already a positive value, so the | | means nothing in this case so Q1 = 1

OpenStudy (jack1):

Q2. 5 |c + b| = 5 | -4 + 3 | =...?